# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v1

@@ Abstract.
We extend the probabilistic approach for constructing Kähler–Einstein metrics on log Fano manifolds $(X,\Delta)$ – involving random point processes – to the case of non-discrete automorphism groups, by explicitly breaking the symmetry. This yields a new algebraic notion of Gibbs polystability, conjecturally equivalent to K-polystability. The definition involves a limit of log canonical thresholds on the GIT semistable locus of the $N$-fold products $X^{N},$ that we conjecture coincides with an analytic reduced stability threshold encoding the coercivity of the K-energy functional modulo automorphisms. These conjectures follow from an overarching conjectural Large Deviation Principle for the limit $N\rightarrow\infty.$ We prove several of our conjectures for log Fano curves and derive a strengthened logarithmic Hardy–Littlewood–Sobolev inequality on the two-sphere, which implies Aubin’s refinement of the Moser–Trudinger inequality. In companion papers we will present applications to Onsager’s point vortex model on the two-sphere and the AdS/CFT correspondence.
@@ 1. Introduction
The quest for canonical metrics on complex manifolds is a classical theme in geometry, tracing back to the uniformization theorem. According to the seminal Yau–Tian–Donaldson conjecture compact complex manifold $X$ admits a Kähler–Einstein metric with positive Ricci curvature, i.e. a Kähler metric with constant positive Ricci curvature, iff $X$ is a K-polystable Fano manifold - this is now a theorem [42, 90]. The Fano condition requires that the anti-canonical line bundle $K_{X}$ of $X$ is ample, while K-polystability is a finer algebro-geometric condition that amounts to a positivity property of all polarized $\mathbb{C}^{*}$–equivariant deformations of $(X,-K_{X}),$ called test configurations. This algebro-geometric characterization has also found striking connections to the AdS/CFT correspondence in theoretical physics, where Kähler–Einstein metrics on Fano manifolds induce Sasaki–Einstein metrics on links that are dual to superconformal gauge theories [82, 44, 43]. More precisely, Fano manifolds yield regular links, while Fano orbifolds give rise to quasi-regular links (and, upon introducing weights, Sasaki–Einstein metrics on general links [6, 76]).
@@ 1.1. Recap of the case when $\text{Aut}_{0}(X)$ is trivial
(1.2) $\nu^{(N)}\coloneqq i^{N^{2}}\big(\mathrm{det}S^{(N)}\big)^{-1/k}\wedge\overline{\left(\det S^{(N)}\right)^{-1/k},}$
where $\det S^{(N)}$ is the holomorphic section of the line bundle $kK_{X^{N}}\rightarrow X^{N},$ defined as the Slater determinant
@@ 1.2. Gibbs polystability and microscopic stability thresholds
We start by introducing a notion of Gibbs polystability, taking the group $\mathcal{G}$ into account. It is formulated in terms of the Zariski open subvariety $(X^{N})_{\text{ss}}$ of $X^{N},$ defined by the semistable locus of $X^{N},$ with respect to the ample $\mathbb{Q}$–line bundle $K_{(X,\Delta)},$ in the sense of Geometric Invariant Theory (GIT) [64]. We will denote by $\Delta_{N}$ the divisor on $X^{N}$ induced by $\Delta$ that is invariant under permutation of the factors of $X^{N}.$ Thus $(X^{N},\Delta_{N})$ is a log Fano manifold. Denote by $\mathcal{D}^{(N)}$ the effective $\mathbb{Q}$-divisor on $X^{N},$ linearly equivalent to $K_{(X^{N},\Delta_{N})},$ cut out by the the $k^{\mathrm{th}}$ root of the section $\det S^{(N)},$ defined by formula 1.3, in terms of a bases in $H^{0}(X,-kK_{(X,\Delta)}).$ To recall the algebraic definition of Gibbs stability, introduced in [13], consider the microscopic stability threshold $\gamma^{(N)}(X,\Delta)$ and its limit $\gamma(X,\Delta)$ defined
@@
(1.7) $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\coloneqq 1+\sup\left\{\epsilon\in\mathbb{R}\colon\exists C_{\epsilon}\in\mathbb{R}\,\,\mathcal{M}\geq\epsilon(\mathcal{I}-\mathcal{J})^{\mathcal{G}}-C_{\epsilon}\right\},$
where $\mathcal{M}$ denotes the (log) Mabuchi functional on the space of Kähler metrics in $c_{1}\left(-K_{(X,\Delta)}\right)$ – whose minimizers are Kähler–Einstein metrics for $(X,\Delta)$ – and the functional $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ denotes the following $\mathcal{G}$–invariant version of the non-negative functional $\mathcal{I}-\mathcal{J},$ defined as the difference of the classical energy functionals $\mathcal{I}$ and $\mathcal{J}$ in Kähler geometry (see Remark 3.1):
(1.8) $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}=\text{inf}_{g\in\mathcal{G}}g^{*}(\mathcal{I}-\mathcal{J}).$
In the case of a complex curve, the three functionals $\mathcal{I}-\mathcal{J},$ $\mathcal{I}$ and $\mathcal{J}$ all coincide with the Dirichlet norm. On a general Kähler manifold they are all comparable and are therefore commonly used as measures of coercivity (or properness) in Kähler geometry [89, 23, 48]; their $\mathcal{G}$–invariant versions appear when studying coercivity modulo the action of $\mathcal{G}$ [95, 48]. In the present context, however, it will be important to work with the functional $\mathcal{I}-\mathcal{J}$ as this choice leads to a thermodynamical interpretation of the corresponding reduced analytic stability threshold $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ (see Prop 3.6).
@@
The validity of this conjecture would imply the second equivalence in Conjecture 1.1. Indeed, by [48], $\delta^{\mathrm{A}}(X,L)^{\mathcal{G}}>1$ iff $(X,\Delta)$ admits a Kähler–Einstein metric, which, by the solution of the YTD conjecture is equivalent to K-polystability (see Theorem 3.7).
@@
The formula $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ implies, in particular, Aubin’s refinement of the Moser–Trudinger inequality on $S^{2}$ [7, 39, 55, 72] and yields a strengthening of the logarithmic Hardy–Littlewood–Sobolev inequality on $S^{2}$ [37, 38] (see Section 4). We show that the discontinuity of $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}},$ as $w$ decreases to $0,$ is a reflection of a somewhat surprising phenomenon of spontaneous symmetry breaking, appearing when $w\in]0,1/2[$ (see Section 4.1). This is related to the fact that the invariant $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}$ then differs from the invariant implicitly appearing in [63, 77], obtained by restricting $\mathcal{M}$ to the subspace of $\mathcal{K}$–invariant Kähler metrics, for a given a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}$ (and replacing $\mathcal{G}$ with its centralizer) – see Section 4.2.
@@ 1.3. Explicit symmetry breaking and the emergence of a unique Kähler–Einstein metric
Given a positive integer $N$ such that 0 is a regular value of $\boldsymbol{m}_{N}$ we denote by $\nu_{\boldsymbol{m}_{N}}^{(N)}$ the measure on $\left\{\boldsymbol{m}_{N}=0\right\}\subset X^{N},$ induced by the measure $\nu^{(N)}$ on $X^{N}$ (defined by formula 1.2 when $\Delta$ is trivial) – see Section 5.1 for the general definition. The algebraic definition of Gibbs polystability of $(X,\Delta)$, introduced in the previous section, is equivalent to the Gibbs semistability of $(X,\Delta)$ together with the analytic assumption that the measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ has finite total mass (Prop 5.11), i.e. that
@@
defines a probability measure on $X^{N},$ supported on $\left\{\boldsymbol{m}_{N}=0\right\}.$ The probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$ is symmetric, i.e. invariant under permutations of $X^{N}$ and $\mathcal{K}$–invariant. The original $\mathcal{G}$–symmetry has thus been reduced to a $\mathcal{K}$–symmetry. Accordingly, the probability measure 1.11 on $X^{N}$ shall be called the the $\mathcal{K}$–reduced canonical probability measure.
We can now view the empirical measure $\delta_{N}$ (formula 1.5) as a random variable on the probability space $\left(X^{N},\mu_{\boldsymbol{m}_{N}}^{(N)}\right).$ Compared to the canonical probability measure 1.4 on $X^{N}$, this means that we are conditioning the $N$ points on $X^{N}$ to have vanishing moment map $\boldsymbol{m}_{N}.$ The empirical measure $\delta_{N}$ of any such configuration of points has vanishing moment, i.e. $\delta_{N}$ maps $X^{N}$ into the space $\mathcal{P}(X)_{0}$ defined in formula 1.11.
@@
where $F$ is the free energy functional, introduced in [10], that may be identified with the Mabuchi functional on the space of all Kähler metrics $\omega$ in $c_{1}(-K_{(X,\Delta)})$ via the relation $F(\omega^{n}/V)=\mathcal{M}(\omega)$. We show that the validity of such a LDP would imply both the asymptotics 1.12 and the convergence 1.13, using that $\mu_{\text{KE}}$ is the unique minimizer of $F(\mu)$ on $\mathcal{P}(X)_{0}$. Moreover, by introducing temperature into the picture the corresponding LDP would also imply Conjecture 1.2, using the observation that the infimum over $\mathcal{G}$ in formula 1.8 is attained at the unique Kähler metric in a given $\mathcal{G}$–orbit whose volume form has vanishing moment (see Prop 3.4).
@@
The convergence 1.12 holds on any K-polystable log Fano curve $(X,\Delta)$ (i.e. when $n=1),$ when $N$ is taken as the sequence of odd positive integers. More precisely,
@@
As for the conjectured convergence towards $\mu_{\text{KE}}$ (formula 1.13) we will show, in the companion paper [4], that it holds up to an arbitrarily small “temperature deformation”, naturally appearing in the context of Onsager’s point vortex model on the two-sphere (see Remark 5.2).
@@
(1.15) $\int_{(X^{N})_{\mathrm{ss}}/\mathcal{G}}\nu^{(N),\mathrm{cc}},$
where $\nu_{\boldsymbol{m}_{N}}^{(N),\mathrm{cc}}$ denotes the $\mathcal{G}$–invariant form on $X^{N}$ of degree $2(N-\text{rank}\mathcal{G})$ obtained by contracting $\nu_{\boldsymbol{m}_{N}}^{(N)}$ with a generator of the top exterior power of the complex Lie algebra $\mathfrak{g}$ of $\mathcal{G}.$ The integral 1.15 is a complex–algebraic period and we show that it can be expressed as a Selberg integral over the field $\mathbb{C}$ [60]. Such integrals can be computed explicitly in terms of Gamma values, as first shown in [53] in the context of Conformal Field Theory, and letting $N\rightarrow\infty$ allows us to deduce Theorem 1.6, by exploiting some connections to K-stable arithmetic log Fano curves in [3].
It should be stressed that, in the general setup introduced above, the sequence of forms $\nu^{(N),\mathrm{cc}}$ appearing in the integral 1.15 is not canonically attached to the given log Fano manifold $(X,\Delta),$ as it depends on the choice of a bases in $H^{0}(X,-kK_{(X,\Delta)})$ and $\mathfrak{g}.$ But, as explained in Section 7, when $(X,\Delta)$ is defined over the integers – or more precisely when $(X,\Delta)$ is the complexification of an arithmetic log Fano variety $(\mathcal{X},\mathcal{D}$) – the arithmetic structure singles out a canonical normalization of $\nu^{(N),\mathrm{cc}}$ and thus a corresponding invariant $\mathcal{Z}_{N}(\mathcal{X},\mathcal{D}).$ Generalizing the K-stable case considered in [3], we conjecture that if $(X,\Delta)$ is K-polystable and the quantized Futaki characters of $(X,\Delta)$ vanish, then $N^{-1}\log\mathcal{Z}_{N}(\mathcal{X},\mathcal{D})$ converges to the infimum of the arithmetic Mabuchi functional of $(\mathcal{X},\mathcal{D}).$ The latter functional, studied in [3], is a logarithmic generalization of Odaka’s modular height [86]. The proof of Theorem 1.6 reveals that this arithmetic conjecture holds for all K-polystable log Fano curves $(X,\Delta)$ defined over the integers.

### v2

@@ Abstract.
We extend the probabilistic approach for constructing Kähler–Einstein metrics on log Fano manifolds $(X,\Delta)$ – involving random point processes – to the case of non-discrete automorphism groups, by explicitly breaking the symmetry. This yields a new algebraic notion of Gibbs polystability, conjecturally equivalent to K-polystability. The definition involves a limit of log canonical thresholds on the GIT semistable locus of the $N$-fold products $X^{N},$ that we conjecture coincides with an analytic reduced stability threshold encoding the coercivity of the K-energy functional modulo automorphisms. These conjectures follow from an overarching conjectural Large Deviation Principle for the limit $N\rightarrow\infty.$ We prove several of our conjectures for log Fano curves. By imposing a moment constraint, we derive a strengthened form of the sharp logarithmic Hardy–Littlewood–Sobolev inequality on the two-sphere, that implies the sharp form of Aubin’s refinement of the Moser–Trudinger inequality. In companion papers we will present applications to Onsager’s point vortex model on the two-sphere and the AdS/CFT correspondence.
@@ 1. Introduction
The quest for canonical metrics on complex manifolds is a classical theme in geometry, tracing back to the uniformization theorem. According to the seminal Yau–Tian–Donaldson conjecture a compact complex manifold $X$ admits a Kähler–Einstein metric with positive Ricci curvature, i.e. a Kähler metric with constant positive Ricci curvature, iff $X$ is a K-polystable Fano manifold - this is now a theorem [42, 90]. The Fano condition requires that the anti-canonical line bundle $K_{X}$ of $X$ is ample, while K-polystability is a finer algebro-geometric condition that amounts to a positivity property of all polarized $\mathbb{C}^{*}$–equivariant deformations of $(X,-K_{X}),$ called test configurations. This algebro-geometric characterization has also found striking connections to the AdS/CFT correspondence in theoretical physics, where Kähler–Einstein metrics on Fano manifolds induce Sasaki–Einstein metrics on links that are dual to superconformal gauge theories [82, 44, 43]. More precisely, Fano manifolds yield regular links, while Fano orbifolds give rise to quasi-regular links (and, upon introducing weights, Sasaki–Einstein metrics on general links [6, 76]).
@@ 1.1. Recap of the case when $\text{Aut}_{0}(X)$ is trivial
(1.2) $\nu^{(N)}\coloneqq\frac{i^{(Nn)^{2}}}{2^{Nn}}\big(\mathrm{det}S^{(N)}\big)^{-1/k}\wedge\overline{\left(\det S^{(N)}\right)^{-1/k},}$
where $n$ is the dimension of $X$ and $\det S^{(N)}$ is the holomorphic section of the line bundle $kK_{X^{N}}\rightarrow X^{N},$ defined as the Slater determinant
@@ 1.2. Gibbs polystability and microscopic stability thresholds
We start by introducing a notion of Gibbs polystability, taking the group $\mathcal{G}$ into account. It is formulated in terms of the Zariski open subvariety $(X^{N})_{\text{ss}}$ of $X^{N},$ defined by the semistable locus of $X^{N},$ with respect to the ample $\mathbb{Q}$–line bundle $K_{(X,\Delta)},$ in the sense of Geometric Invariant Theory (GIT) [64]. We will denote by $\Delta_{N}$ the divisor on $X^{N}$ induced by $\Delta$ that is invariant under permutation of the factors of $X^{N}.$ Thus $(X^{N},\Delta_{N})$ is a log Fano manifold. Denote by $\mathcal{D}^{(N)}$ the effective $\mathbb{Q}$-divisor on $X^{N},$ linearly equivalent to $K_{(X^{N},\Delta_{N})},$ cut out by the the $k^{\mathrm{th}}$ root of the section $\det S^{(N)},$ defined by formula 1.3, in terms of a bases in $H^{0}(X,-kK_{(X,\Delta)}).$ To recall the algebraic definition of Gibbs stability, introduced in [13], consider the microscopic stability threshold $\gamma^{(N)}(X,\Delta)$ and its limit $\gamma(X,\Delta)$ defined by
@@
(1.7) $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\coloneqq 1+\sup\left\{\epsilon\in\mathbb{R}\colon\exists C_{\epsilon}\in\mathbb{R}\,\,\mathcal{M}\geq\epsilon(\mathcal{I}-\mathcal{J})^{\mathcal{G}}+C_{\epsilon}\right\},$
where $\mathcal{M}$ denotes the (log) Mabuchi functional on the space of Kähler metrics in $c_{1}\left(-K_{(X,\Delta)}\right),$ whose minimizers are Kähler–Einstein metrics for $(X,\Delta),$ and $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ denotes the following $\mathcal{G}$–invariant version of the non-negative functional $\mathcal{I}-\mathcal{J},$ defined as the difference of the classical energy functionals $\mathcal{I}$ and $\mathcal{J}$ in Kähler geometry
(1.8) $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}=\text{inf}_{g\in\mathcal{G}}g^{*}(\mathcal{I}-\mathcal{J})$
(the definition of $\mathcal{I}$ and $\mathcal{J}$ are recalled in Remark 3.1). In the case of a complex curve, the three functionals $\mathcal{I}-\mathcal{J},$ $\mathcal{I}$ and $\mathcal{J}$ all coincide with the square of the Dirichlet norm of the Kähler potential. On a general Kähler manifold they are all comparable and are therefore commonly used as measures of coercivity (or properness) in Kähler geometry [89, 23, 48]; their $\mathcal{G}$–invariant versions appear when studying coercivity modulo the action of $\mathcal{G}$ [95, 48]. In the present context, however, it will be important to work with the functional $\mathcal{I}-\mathcal{J}$ as this choice leads to a thermodynamical interpretation of the corresponding reduced analytic stability threshold $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ (see Prop 3.6).
@@
The validity of this conjecture would imply the second equivalence in Conjecture 1.1. Indeed, as shown in [48], confirming a conjecture of Tian, $\delta^{\mathrm{A}}(X,L)^{\mathcal{G}}>1$ iff $(X,\Delta)$ admits a Kähler–Einstein metric, which, by the solution of the YTD conjecture is equivalent to K-polystability (see Theorem 3.7).
@@
The formula $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ implies, in particular, Aubin’s refinement of the Moser–Trudinger inequality on $S^{2}$ [7, 39, 55, 72] and yields a strengthening of the sharp logarithmic Hardy–Littlewood–Sobolev inequality on $S^{2}$ in [37, 38, 9] under a moment constraint, i.e. a strengthening of the following inequality for all probability measures $\mu$ on $S^{2}:$
(1.9) $E(\mu)\leq D(\mu|\mu_{0}),$
where $D(\mu|\mu_{0})$ denotes the entropy of $\mu,$ relative to the $SO(3)$-invariant probability measure $\mu_{0}$ on $S^{2},$ and $E(\mu)$ denotes the logarithmic energy of $\mu$ (explicitely expressed in formula 4.3). In fact, combining our formula for $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}$ with the uniqueness result for the Liouville equation on $S^{2}$ established in [72] yields:
Theorem 1.4.
Let $X=\mathbb{P}^{1},$ identified with the unit-sphere $S^{2}$ in $\mathbb{R}^{3},$ and denote by $\mathcal{M}$ the Mabuchi functional on the space of all Kähler forms $\omega$ in $c_{1}(X),$ normalized so that $\mathcal{M}(\omega)=0$ when $\omega$ is proportional to the $SO(3)$-invariant probability measure $\mu_{0}$ on $S^{2}.$ For any $\omega$ in $c_{1}(X),$
$\mathcal{M}(\omega)\geq(\mathcal{I}-\mathcal{J})^{\mathcal{G}}(\omega)$
with equality iff $\omega$ is proportional to $\mu_{0}$ modulo the action of $\mathcal{G}$ (i.e. modulo conformal diffeomorphisms of $S^{2}).$ Moreover, the functional $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ may be replaced by $\mathcal{I}-\mathcal{J}$ if the Kähler potential of $\omega$ satisfies Tian’s orthogonality condition 3.31. More generally, for any probability measure $\mu$ on $S^{2},$ satisfying the moment constraint $\int_{S^{2}\subset\mathbb{R}^{3}}\boldsymbol{x}\mu=0,$
(1.10) $E(\mu)\leq\frac{1}{2}D(\mu|\mu_{0})$
with equality iff $\mu=\mu_{0}.$
The inequality for $\mathcal{M}$ in formula 1.7, for some (non-effective) positive constant $\epsilon$ and $C_{\epsilon}\in\mathbb{R}$ was first established in [48]. Since $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ (by Theorem 1.3) the inequality for $\mathcal{M}$ fails on $\mathbb{P}^{1}$ when $\epsilon>1$ (for any $C_{\epsilon})$ and thus the multiplicative constants in Theorem 1.4 are sharp. More generally, the formula for $\delta^{\mathrm{A}}(X,w)^{\mathcal{G}}$ in Theorem 1.3 yields a quasi-sharp generalization of Aubins’ improvement of the Moser-Trudinger inequality to $w\neq 0.$ We show that the discontinuity of $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}},$ as $w$ decreases to $0,$ is a reflection of a somewhat surprising phenomenon of spontaneous symmetry breaking, appearing when $w\in]0,1/2[$ (see Section 4.1). This is related to the fact that the invariant $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}$ then differs from the invariant obtained by restricting $\mathcal{M}$ to the subspace of $\mathcal{K}$–invariant Kähler metrics, for a given a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}$ (and replacing $\mathcal{G}$ with its center), appearing implicitly in [48, 63, 77] – see Section 4.2.
@@ 1.3. Explicit symmetry breaking and the emergence of a unique Kähler–Einstein metric
Given a positive integer $N$ such that 0 is a regular value of $\boldsymbol{m}_{N}$ we denote by $\nu_{\boldsymbol{m}_{N}}^{(N)}$ the measure on $\left\{\boldsymbol{m}_{N}=0\right\}\subset X^{N},$ induced by the measure $\nu^{(N)}$ on $X^{N}$ (defined by formula 1.2 when $\Delta$ is trivial) – see Section 6.1 for the general definition. The algebraic definition of Gibbs polystability of $(X,\Delta)$, introduced in the previous section, is equivalent to the Gibbs semistability of $(X,\Delta)$ together with the analytic assumption that the measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ has finite total mass (Prop 6.11), i.e. that
@@
defines a probability measure on $X^{N},$ supported on $\left\{\boldsymbol{m}_{N}=0\right\}.$ The probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$ is symmetric, i.e. invariant under permutations of $X^{N}$ and $\mathcal{K}$–invariant. The original $\mathcal{G}$–symmetry has thus been reduced to a $\mathcal{K}$–symmetry. Accordingly, the probability measure 1.13 on $X^{N}$ shall be called the the $\mathcal{K}$–reduced canonical probability measure.
We can now view the empirical measure $\delta_{N}$ (formula 1.5) as a random variable on the probability space $\left(X^{N},\mu_{\boldsymbol{m}_{N}}^{(N)}\right).$ Compared to the canonical probability measure 1.4 on $X^{N}$, this means that we are conditioning the $N$ points on $X^{N}$ to have vanishing moment map $\boldsymbol{m}_{N}.$ The empirical measure $\delta_{N}$ of any such configuration of points has vanishing moment, i.e. $\delta_{N}$ maps $X^{N}$ into the space $\mathcal{P}(X)_{0}$ defined in formula 1.13.
@@
where $F$ is the free energy functional, introduced in [10], that may be identified with the Mabuchi functional on the space of all Kähler metrics $\omega$ in $c_{1}(-K_{(X,\Delta)})$ via the relation $F(\omega^{n}/V)=\mathcal{M}(\omega)$. We show that the validity of such a LDP would imply both the asymptotics 1.14 and the convergence 1.15, using that $\mu_{\text{KE}}$ is the unique minimizer of $F(\mu)$ on $\mathcal{P}(X)_{0}$. Moreover, by introducing temperature into the picture the corresponding LDP would also imply Conjecture 1.2, using the observation that the infimum over $\mathcal{G}$ in formula 1.8 is attained at the unique Kähler metric in a given $\mathcal{G}$–orbit whose volume form has vanishing moment (see Prop 3.4).
@@
The convergence 1.14 holds on any K-polystable log Fano curve $(X,\Delta)$ (i.e. when $n=1$), when $N$ is taken as the sequence of odd positive integers. More precisely,
@@
As for the conjectured convergence towards $\mu_{\text{KE}}$ (formula 1.15) we will show, in the companion paper [4], that it holds up to an arbitrarily small “temperature deformation”, naturally appearing in the context of Onsager’s point vortex model on the two-sphere (see Remark 6.2).
@@
(1.17) $\int_{(X^{N})_{\mathrm{ss}}/\mathcal{G}}\nu_{\mathbf{m}_{N}}^{(N),\mathrm{cc}},$
where $\nu_{\boldsymbol{m}_{N}}^{(N),\mathrm{cc}}$ denotes the $\mathcal{G}$–invariant form on $X^{N}$ of degree $2(N-\text{rank}\mathcal{G})$ obtained by contracting $\nu_{\boldsymbol{m}_{N}}^{(N)}$ with a generator of the top exterior power of the complex Lie algebra $\mathfrak{g}$ of $\mathcal{G}.$ The integral 1.17 is a complex–algebraic period and we show that it can be expressed as a Selberg integral over the field $\mathbb{C}$ [60]. Such integrals can be computed explicitly in terms of Gamma values, as first shown in [53] in the context of Conformal Field Theory, and letting $N\rightarrow\infty$ allows us to deduce Theorem 1.7, by exploiting some connections to K-stable arithmetic log Fano curves in [3].
It should be stressed that, in the general setup introduced above, the sequence of forms $\nu^{(N),\mathrm{cc}}$ appearing in the integral 1.17 is not canonically attached to the given log Fano manifold $(X,\Delta),$ as it depends on the choice of a bases in $H^{0}(X,-kK_{(X,\Delta)})$ and $\mathfrak{g}.$ But, as explained in Section 8, when $(X,\Delta)$ is defined over the integers – or more precisely when $(X,\Delta)$ is the complexification of an arithmetic log Fano variety $(\mathcal{X},\mathcal{D}$) – the arithmetic structure singles out a canonical normalization of $\nu^{(N),\mathrm{cc}}$ and thus a corresponding invariant $\mathcal{Z}_{N}(\mathcal{X},\mathcal{D}).$ Generalizing the K-stable case considered in [3], we conjecture that if $(X,\Delta)$ is K-polystable and the quantized Futaki characters of $(X,\Delta)$ vanish, then $N^{-1}\log\mathcal{Z}_{N}(\mathcal{X},\mathcal{D})$ converges to the infimum of the arithmetic Mabuchi functional of $(\mathcal{X},\mathcal{D}).$ The latter functional, studied in [3], is a logarithmic generalization of Odaka’s modular height [86]. The proof of Theorem 1.7 reveals that this arithmetic conjecture holds for all K-polystable log Fano curves $(X,\Delta)$ defined over the integers.